OpenStudy (lgbasallote):

scientific studies show that DNA deteriorates over time. if excavations and diggings of fossils indicate that only 1.2% of DNA survive after 25 million years, what percent of an original sample would have been lost after 10 million years?

OpenStudy (lgbasallote):

apply exponential decay formula \[\huge x=ce^{−kt}\] isolate t \[\huge \ln(\frac xc)= −kt\] \[\huge \implies −\frac{\ln(\frac xc)}k=t\] simplify \[\huge \frac{ln(\frac cx)}k=t\] after 2.5 M years, 1.2% of DNA survived let c be the original value. 1.2% is 0.012c \[\huge \frac {\ln(\frac c{0.012c})}k=2.5\] solve for k \[\huge \implies k=\frac{\ln(\frac c{0.012c})}{2.5}\] simplify \[\huge k= \frac{\ln(83.3\overline 3)}{2.5}\] solve for k \[\huge \implies k=1.7691\]

OpenStudy (lgbasallote):

now...solve for the percentage remaining after 10 million years \[\huge \implies \frac{\ln (\frac c{ac})}{0.17691} = 10\] where: a = percentage now...solve for a \[\huge \implies \ln (\frac c{ac}) = 1.7691\] \[\huge \frac c{ac} = 5.8656\] \[\huge \frac 1a = 5.8656\] \[\huge \frac 1{5.8656} = a\] \[\huge a = 0.1705\] this is the amount remaining. to solve the amount lost, subtract that from 1 \[\huge \implies 1 - 0.1705 = 0.8295\] so.. 82.95% was lost

OpenStudy (lgbasallote):

got it @Jlastino ?

OpenStudy (jlastino):

yep...thanks!